# Symbolic Feedforward Networks for Probabilistic Finite Automata: Exact Simulation and Learnability

### v1

@@ 7 Conclusion
This paper presented a formal, constructive, and learnable framework for simulating probabilistic finite automata (PFAs) using symbolic feedforward neural networks. By extending recent advances in automata-as-networks theory to the probabilistic domain, we established a suite of theoretical results showing that feedforward networks with matrix-vector updates and soft activations can precisely simulate PFA behavior, including probabilistic subset construction, $\varepsilon$-closure, and symbolic state propagation. Our construction preserves probabilistic semantics without requiring recurrence, sampling, or external memory, and supports exact simulation, differentiable learning, and parameter efficiency.
We further demonstrated that such symbolic PFA simulators can not only encode PFAs but also learn them directly from labeled sequences data using standard gradient descent-based optimization, achieving perfect generalization and convergence across seeds. This provides strong evidence that symbolic neural architectures are capable not only of encoding formal probabilistic models, but also of learning them directly from labeled sequences. Our results advance the frontier of neural-symbolic computation, unifying classical automata theory with modern deep learning tools under a rigorous probabilistic framework.
@@ Appendix A Appendix
$\sum_{j=1}^{n}[s_{t+1}]_{j}=\sum_{j=1}^{n}\sum_{i=1}^{n}[s_{t}]_{i}\cdot T^{x_{t}}_{ij}=\sum_{i=1}^{n}[s_{t}]_{i}\left(\sum_{j=1}^{n}T^{x_{t}}_{ij}\right)=\sum_{i=1}^{n}[s_{t}]_{i}\cdot 1=1,$
@@
$\bm{p}^{(k+1)}=\sigma\!\left(\bm{p}^{(k)}T^{\varepsilon}\right),\quad\bm{p}^{(0)}=\bm{p}_{t},$
where $\sigma$ acts elementwise as $\sigma(z)=\min(z,1)$ to ensure all values lie within $[0,1]$ and prevent probability overflow. Each update propagates $\varepsilon$-transitions from the current distribution.
Monotonicity: Since $T^{\varepsilon}$ has nonnegative entries and $\sigma$ is monotonic, the sequence $\bm{p}^{(k)}$ is elementwise non-decreasing:
@@
Boundedness: Because $\bm{p}^{(0)}\in[0,1]^{1\times n}$ and $\sigma$ clips each entry to $\leq 1$, the sequence is bounded: $\bm{p}^{(k)}\leq\mathbf{1}$ for all $k$.
@@
Proof.
Let the initial state vector be $s_{0}:=\pi_{0}\in\Delta^{n}$. For each symbol $x_{t}$ in the input sequence $x=x_{1}\cdots x_{L}$, define a symbolic linear transformation by the matrix $T^{x_{t}}\in\mathcal{S}_{n}$. Then define the intermediate vectors via recurrence:
$s_{t}=s_{t-1}T^{x_{t}},\quad\text{for }t=1,\dots,L.$
By Proposition 4.1 and Theorem 4.5, each $s_{t}\in\Delta^{n}$ remains a valid probabilistic state vector and encodes the marginal distribution over states after consuming the first $t$ symbols.
Finally, define the network output as the inner product:
$f_{\theta}(x)=\langle s_{L},\mathbf{1}_{F}\rangle=\sum_{i\in F}[s_{L}]_{i}.$
This equals the total probability mass assigned to accepting states after processing the full input, which matches the formal definition of $P_{\mathcal{A}}(x)$.
Since each operation in this computation is a linear transformation with fixed symbolic matrices followed by no activation (or identity), the computation for any length-$L$ input induces a depth-$L$ feedforward computation graph of width $n$ (one layer per symbol), with shared matrix blocks per symbol. ∎
A.5 Proof of Proposition  4.9: Parameter Efficiency of Feedforward Automata Simulators
Proof.
Each symbol $x\in\Sigma$ is associated with a transition matrix $T^{x}\in\mathbb{R}^{n\times n}$, where $T^{x}_{ij}\in[0,1]$ and $\sum_{j}T^{x}_{ij}=1$. Each matrix contains $n^{2}$ real-valued entries, which may either be stored directly or parameterized via softmax normalization over rows.
Thus, storing all $k$ matrices requires $k\cdot n^{2}$ real parameters. The only other components of the model are:
@@
Hence, the total number of trainable parameters is $O(kn^{2})$, regardless of the input length $L$, since the same transition matrices are reused at each layer of the feedforward network. ∎
@@
Forward Direction: Let $\mathcal{A}=(Q,\Sigma,\delta,q_{0},F)$ be a PFA with $|Q|=n$. For each symbol $x\in\Sigma$, let $T^{x}\in[0,1]^{n\times n}$ be its transition matrix, where each row $i$ defines a distribution $\sum_{j}(T^{x})_{ij}=1$ and $(T^{x})_{ij}=P(q_{i}\to q_{j}\mid x)$.
Initialization: Let $s_{0}\in\mathbb{R}^{1\times n}$ be the one-hot row vector with $s_{0}]_{i}=\mathbf{1}[q_{i}=q_{0}]$.
@@
$s_{t}=s_{t-1}T^{x_{t}}.$
Each $s_{t}\in\mathbb{R}^{1\times n}$ is a probability distribution over states after reading $x_{1}\dots x_{t}$. This follows because $s_{0}$ is one-hot and each $T^{x_{t}}$ is row-stochastic, preserving the $L_{1}$ norm.
@@
which is precisely the total probability that the PFA reaches any accepting state after reading $x$.
@@
  • The state vector is updated via $s_{t}=s_{t-1}T^{x_{t}}$,
@@
We now define a PFA $\mathcal{A}^{\prime}=(Q,\Sigma,\delta^{\prime},q_{0},F)$ such that:
@@
  • For every $x\in\Sigma$, define $\delta^{\prime}(q_{i},x)(q_{j})=(T^{x})_{ij}$.
Because each $T^{x}$ is row-stochastic by construction, this defines a valid probabilistic transition function. The forward recursion of this PFA will match the feedforward network’s propagation, and the acceptance probability is computed in the same way via $\langle s_{L},1_{F}\rangle$.
@@
Proof.
Let $\mathcal{A}=(Q,\Sigma,\delta,q_{0},F)$ be a PFA with $|Q|=n$ states. We construct a symbolic feedforward network $f_{\theta}$ as follows:
1. Input and Transition Representation. Each input symbol $x\in\Sigma$ is associated with a stochastic transition matrix $T^{x}\in[0,1]^{n\times n}$, satisfying $\sum_{j=1}^{n}T^{x}_{ij}=1$ for all $i$ (row-stochastic). These matrices are initialized randomly with normalization and treated as trainable parameters.
2. Initial State. Let $s_{0}=e_{q_{0}}\in\mathbb{R}^{1\times n}$ be the one-hot row vector encoding the initial state.
@@
$s_{t}=s_{t-1}T^{x_{t}},\quad\text{for }t=1,2,\dots,L.$
@@
5. Training Objective. Given a labeled dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{m}$ with ground-truth acceptance probabilities $y_{i}=P_{\mathcal{A}}(x_{i})$, we train the network using mean squared error (MSE) loss:
$\mathcal{L}(\theta)=\frac{1}{m}\sum_{i=1}^{m}\left(f_{\theta}(x_{i})-y_{i}\right)^{2}.$
6. Optimization. The transition matrices $\{T^{x}\}_{x\in\Sigma}$ are optimized using gradient-based updates:
@@
with optional projection after each update to ensure each $T^{x}$ remains row-stochastic (e.g., row-wise softmax or Sinkhorn-style normalization).
7. Empirical Approximation. Because the model class includes exact simulators (Theorem 4.10) and the training objective directly penalizes deviations from labels, the empirical risk can be driven arbitrarily low on $\mathcal{D}$ given sufficient optimization and capacity. In particular, when labels are realizable by some $\{T^{x}\}$, there exist parameters with zero training error. ∎

### v2

@@ 7 Conclusion
This paper presented a formal, constructive, and learnable framework for simulating probabilistic finite automata (PFAs) with $\varepsilon$-transitions using symbolic feedforward neural networks. By extending recent advances in automata-as-networks theory to the probabilistic domain, we established a suite of theoretical results showing that feedforward networks with matrix-vector updates and soft activations can precisely simulate PFA behavior, including probabilistic subset construction, finite $\varepsilon$-closure and symbolic state propagation. Our construction preserves probabilistic semantics and supports exact simulation, differentiable learning, and parameter efficiency.
We further demonstrated that such symbolic PFA simulators can not only encode probabilistic automata but also learn them directly from labeled sequence data using standard gradient descent-based optimization, achieving perfect generalization and convergence across seeds. This provides strong evidence that symbolic neural architectures are capable not only of encoding formal probabilistic models, but also of learning them directly from labeled data. Our results advance the frontier of neural-symbolic computation, unifying classical automata theory with modern deep learning tools under a rigorous probabilistic framework.
@@ Appendix A Appendix
$\sum_{j=1}^{n}[s_{t+1}]_{j}$ $=\sum_{j=1}^{n}\sum_{i=1}^{n}[s_{t}]_{i}\cdot T^{x_{t}}_{ij}$ (2)
$=\sum_{i=1}^{n}[s_{t}]_{i}\left(\sum_{j=1}^{n}T^{x_{t}}_{ij}\right)$ (3)
$=\sum_{i=1}^{n}[s_{t}]_{i}\cdot 1$ (4)
$=1$ (5)
@@
$\bm{p}^{(k+1)}=\bm{p}^{(k)}T^{\varepsilon},\quad\bm{p}^{(0)}=\bm{p}_{t}.$
Monotonicity: Since $T^{\varepsilon}$ has nonnegative entries, the sequence $\bm{p}^{(k)}$ is elementwise non-decreasing:
@@
Boundedness: Because $\bm{p}^{(0)}\in[0,1]^{1\times n}$ and $T^{\varepsilon}$ is row-substochastic, the sequence remains bounded in $0,1]^{1\times n}$.
@@
Proof sketch.
Let the initial state vector be $s_{0}:=\pi_{0}\in\Delta^{n}$, and apply the finite $\varepsilon$-closure:
$s_{0}:=\pi_{0}\Big(\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}\Big).$
For each symbol $x_{t}$ in the input sequence $x=x_{1}\cdots x_{L}$, define the recurrence
$s_{t}=s_{t-1}\,T^{x_{t}}\Big(\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}\Big),\quad\text{for }t=1,\dots,L.$
By Proposition 4.1, Lemma 4.6, and Theorem 4.5, each $s_{t}\in\Delta^{n}$ remains a valid probabilistic state vector encoding the distribution after consuming the first $t$ symbols and all intervening $\varepsilon$-steps. Finally, define the network output as
$f_{\theta}(x)=\langle s_{L},\mathbf{1}_{F}\rangle=\sum_{i\in F}[s_{L}]_{i},$
which equals the acceptance probability of $\mathcal{A}$ on $x$. Since each operation is a symbolic linear transformation with finite $\varepsilon$-closure, the computation for any length-$L$ input induces a depth-$L$ feedforward computation graph of width $n$. ∎
A.5 Proof of Proposition 4.9: Parameter Efficiency of Feedforward Automata Simulators
Proof sketch.
Each symbol $x\in\Sigma$ is associated with a transition matrix $T^{x}\in\mathbb{R}^{n\times n}$, where $T^{x}_{ij}\in[0,1]$ and $\sum_{j}T^{x}_{ij}=1$. Each such matrix contains $n^{2}$ real-valued entries, which may either be stored directly or parameterized via row-wise softmax normalization. In addition, there is a single $\varepsilon$-transition matrix $T^{\varepsilon}\in\mathbb{R}^{n\times n}$, which also contributes $n^{2}$ parameters. The finite closure operator $(T^{\varepsilon})^{*}=\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}$ is computed symbolically and introduces no additional parameters.
Thus, storing all $k$ symbol matrices plus the $\varepsilon$-matrix requires $(k+1)\cdot n^{2}$ real parameters. The only other components of the model are:
@@
Hence, the total number of trainable parameters is $O(kn^{2})$, regardless of the input length $L$, since the same transition matrices (including $T^{\varepsilon}$) are reused at each layer of the feedforward network. ∎
@@
Forward Direction: Let $\mathcal{A}=(Q,\Sigma\cup\{\varepsilon\},\delta,q_{0},F)$ be a PFA with $|Q|=n$. For each symbol $x\in\Sigma$, let $T^{x}\in[0,1]^{n\times n}$ be its transition matrix, where each row $i$ defines a distribution $\sum_{j}(T^{x})_{ij}=1$. Let $T^{\varepsilon}\in[0,1]^{n\times n}$ denote the $\varepsilon$-transition matrix, and define the finite closure operator $(T^{\varepsilon})^{*}=\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}$.
Initialization: Let $s_{0}\in\mathbb{R}^{1\times n}$ be the one-hot row vector with $s_{0}]_{i}=\mathbf{1}[q_{i}=q_{0}]$, and apply closure:
$s_{0}=s_{0}(T^{\varepsilon})^{*}.$
@@
$s_{t}=s_{t-1}T^{x_{t}}(T^{\varepsilon})^{*}.$
Each $s_{t}\in\mathbb{R}^{1\times n}$ is a probability distribution over states after reading $x_{1}\dots x_{t}$ and applying all intervening $\varepsilon$-steps. This follows because $s_{0}$ is one-hot, each $T^{x_{t}}$ is row-stochastic, and $(T^{\varepsilon})^{*}$ preserves normalization.
@@
which is precisely the acceptance probability of the PFA.
@@
  • An $\varepsilon$-transition matrix $T^{\varepsilon}$ is applied via the finite closure operator $(T^{\varepsilon})^{*}$,
  • The state vector is updated via $s_{t}=s_{t-1}T^{x_{t}}(T^{\varepsilon})^{*}$,
@@
We now define a PFA $\mathcal{A}^{\prime}=(Q,\Sigma\cup\{\varepsilon\},\delta^{\prime},q_{0},F)$ such that:
@@
  • For every $x\in\Sigma$, define $\delta^{\prime}(q_{i},x)(q_{j})=(T^{x})_{ij}$,
  • For the $\varepsilon$-transition, define $\delta^{\prime}(q_{i},\varepsilon)(q_{j})=(T^{\varepsilon})_{ij}$.
Because each $T^{x}$ and $T^{\varepsilon}$ is row-stochastic by construction, this defines a valid probabilistic transition function with finite closure. The forward recursion of this PFA matches the feedforward network’s propagation, and the acceptance probability is computed in the same way via $\langle s_{L},1_{F}\rangle$.
@@
Proof sketch.
Let $\mathcal{A}=(Q,\Sigma\cup\{\varepsilon\},\delta,q_{0},F)$ be a PFA with $|Q|=n$ states. We construct a symbolic feedforward network $f_{\theta}$ as follows:
1. Input and Transition Representation. Each input symbol $x\in\Sigma$ is associated with a stochastic transition matrix $T^{x}\in[0,1]^{n\times n}$, satisfying $\sum_{j=1}^{n}T^{x}_{ij}=1$ for all $i$ (row-stochastic). In addition, there is a single $\varepsilon$-transition matrix $T^{\varepsilon}\in[0,1]^{n\times n}$. These matrices are initialized randomly with normalization and treated as trainable parameters.
2. Initial State. Let $s_{0}=e_{q_{0}}\in\mathbb{R}^{1\times n}$ be the one-hot row vector encoding the initial state. Apply the finite $\varepsilon$-closure:
$s_{0}=s_{0}(T^{\varepsilon})^{*},\quad(T^{\varepsilon})^{*}=\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}.$
@@
$s_{t}=s_{t-1}T^{x_{t}}(T^{\varepsilon})^{*},\quad\text{for }t=1,2,\dots,L.$
@@
5. Training Objective. Given a labeled dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{m}$ with ground-truth acceptance probabilities $y_{i}=P_{\mathcal{A}}(x_{i})$, we train the network using binary cross-entropy (BCE) loss:
$\mathcal{L}(\theta)=-\frac{1}{m}\sum_{i=1}^{m}\Big(y_{i}\log f_{\theta}(x_{i})+(1-y_{i})\log(1-f_{\theta}(x_{i}))\Big).$
6. Optimization. The transition matrices $\{T^{x}\}_{x\in\Sigma}$ and $T^{\varepsilon}$ are optimized using gradient-based updates:
@@
with optional projection after each update to ensure row-stochasticity (e.g., row-wise softmax or Sinkhorn-style normalization).
7. Empirical Approximation. Because the model class includes exact simulators (Theorem 4.10) and the training objective directly penalizes deviations from labels, the empirical risk can be driven arbitrarily low on $\mathcal{D}$ given sufficient optimization and capacity. In particular, when labels are realizable by some $\{T^{x},T^{\varepsilon}\}$, there exist parameters with zero training error. ∎